OpenStudy (love_to_love_you):

Will medal and fan Simplify 3-2(x+4). Show all work.

OpenStudy (alexandervonhumboldt2):

3-2x-8=-5-2x

OpenStudy (love_to_love_you):

wow that simple?

OpenStudy (alexandervonhumboldt2):

yeash

OpenStudy (love_to_love_you):

Could you help me with another?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (love_to_love_you):

Solve this multiple step equation. Show all work. (m-2)-5=8-2(m-4)

OpenStudy (alexandervonhumboldt2):

m-2-5=8-2m+8

OpenStudy (alexandervonhumboldt2):

3m=16+2+5

OpenStudy (alexandervonhumboldt2):

3m=23

OpenStudy (alexandervonhumboldt2):

m=23/3

OpenStudy (love_to_love_you):

thank youuuu

OpenStudy (love_to_love_you):

1 more? haha sorry!

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

post!

OpenStudy (love_to_love_you):

okay

OpenStudy (love_to_love_you):

Solve and graph each compound inequality. Show all work. a) 3x ≥ -12 and 8x ≤ 16

OpenStudy (love_to_love_you):

Solve and graph each compound inequality. Show all work. a) 3x >= -12 and 8x <= 16

OpenStudy (love_to_love_you):

b) 3x >= 3 or 9x < 54

OpenStudy (alexandervonhumboldt2):

A) x>=-4 x<=2

OpenStudy (alexandervonhumboldt2):

b) x>=1 x<6

OpenStudy (love_to_love_you):

how do I graph that?

OpenStudy (alexandervonhumboldt2):

https://www.mathway.com/graph

OpenStudy (alexandervonhumboldt2):

enter the simplified form i gave you

OpenStudy (love_to_love_you):

ok thanks

OpenStudy (alexandervonhumboldt2):

tag me aytime but try to post each question in separate post

OpenStudy (love_to_love_you):

Alrighty, will do. Thank you sooooo much!

OpenStudy (alexandervonhumboldt2):

ok!